OpenStudy (anonymous):

Kylie recorded the number of pears on a pear tree at the end of every week. The following function represents the number of pears on the pear tree, where x represents the number of weeks. P(x) = 2^(2 + x) Which statement is true? The number of pears on the tree quadruples each week. The number of pears on the tree increases by 2 each week. The number of pears on the tree increases by 4 each week. The number of pears on the tree doubles each week.

OpenStudy (anonymous):

nvmd i see that it would be D First, use the properties of exponents to rewrite the function. P(x) = 2^(2 + x) = 2^2 • 2^x = 4 • 2^x = 4(1 + 1)^x In the expression, 4(1 + 1)x, four is the initial number of pears on the tree, and (1 + 1)x reveals the rate of increase (100%) of the number of pears on the tree. Therefore, the number of pears on the tree doubles each week.